OpenStudy (wilkes688):

A parabola is defined by the equation (x − 5)2 = 12(y + 2). In which direction will the parabola open?

OpenStudy (evoker):

You need to look at the squared part when the non squared part has no coefficient in front of it.

OpenStudy (evoker):

basically if you have y=+x^2 opens up, y=-x^2 opens down, x=+y^2 opens right and x=-y^2 opens left.